OpenStudy (anonymous):

Sal's Sandwich Shop sells wraps and sandwiches as part of its lunch specials. The profit on every sandwich is $2 and the profit on every wrap is $3. Sal made a profit of $1,470 from lunch specials last month. The equation 2x + 3y = 1,470 represents Sal's profits last month, where x is the number of sandwich lunch specials sold and y is the number of wrap lunch specials sold. Change the equation to slope-intercept form. Identify the slope and y-intercept of the equation. Be sure to show all your work Describe how you would graph this line using the slope-intercept methodBe sure to write usi

OpenStudy (anonymous):

\[2x+3y=1470\] The first thing you need to do is rearrange the formula so only the Y is on the left hand side .All you need to do is do the opposite to what is given

OpenStudy (anonymous):

So in this case you would start by subtracting 2x from the left hand side so ,so you hav eto minus 2x from both sides It should look like this \[3y = -2x + 1470\]

OpenStudy (anonymous):

Then you still have 3y instead of y so you need to divide both sides by 3 . So you need to divide 3x by 3,-2x by 3 and 1470 by 3 It should look like this \[\frac{ 3y }{ 3}=\frac{ -2x }{ 3 }+\frac{ 1470 }{ 3 }\]

OpenStudy (anonymous):

The equation should look like this\[y= -\frac{ 2 }{ 3 }x +490\]

OpenStudy (anonymous):

Now to graph this line you can find the x and y intercepts Let start with the y intercept To do this all you need to do is substitute the x for 0 to find y

OpenStudy (anonymous):

y int x=0 \[y=(-\frac{ 2 }{ 3 }\times0)+490\] y=490

OpenStudy (anonymous):

x intercept can be found when y = 0 \[0=-\frac{ 2 }{ 3}x+ 490\] Then solve it like a linear equation So 0-490 and then divide this by -2/3 It shoudld equal 735

OpenStudy (anonymous):

To graph this so can see that you need a big scale because of the numbers so a graph like the one below should suffice

OpenStudy (anonymous):

The scale for the graph should be 1cm for 100 and then just plot the points On the y axis,plot it at 490 and on the x axis ,plot it at 735 and then draw a line between the two

OpenStudy (anonymous):



OpenStudy (anonymous):

Hope that helps @ripnana813